OpenStudy (anonymous):

Simplify √24/18 (simplify square root of 24 over 18)?

OpenStudy (anonymous):

=sqrt(4/3)=2sqrt(1/3)

OpenStudy (solomonzelman):

It looks like this right? \(\large\color{midnightblue}{ \rm \sqrt{ \frac{24}{18} } }\)

OpenStudy (anonymous):

correct

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm \sqrt{ \frac{24}{18} } }\) gives, \(\large\color{midnightblue}{ \rm \frac{ \sqrt{24} }{ \sqrt{18} } }\) \(\large\color{midnightblue}{ \rm \frac{ 2\sqrt{6} }{ 3\sqrt{2} } }\) \(\large\color{midnightblue}{ \rm \frac{ 2\sqrt{6} ~~~~\times \sqrt{2}}{ 3\sqrt{2} ~~~~\times \sqrt{2} } }\)

OpenStudy (solomonzelman):

can you finish the problem ?

OpenStudy (anonymous):

2 square root 12/ 3 square root 6?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Close:) \(\huge\color{blue}{ \frac{ 2 \sqrt{12} }{ 3 \sqrt{4} } }\)

OpenStudy (solomonzelman):

and 4 is a perfect square. Juts simplify top and bottom:)

OpenStudy (anonymous):

4 square root 3/6?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

nvm I got it

OpenStudy (solomonzelman):

yes, and 4 and 6 are both divisible by 2, so the final answer is ?

OpenStudy (solomonzelman):

1) Click and Hold ALT 2) Click 2 5 1 (on the keypad on the right, and not the numbers below F1 F2 F3...) 3) release the ALT, •••• you get √

OpenStudy (anonymous):

Thanks, but my keyboard does not have numbers on the right

OpenStudy (solomonzelman):

you don't have the number pad on the right of the keyboard?? unusual, I must say :) -:(

OpenStudy (solomonzelman):

Anytime ...